Th combined SAT scores for students taking the SAT-I tests are normally distributed with a mean of 982 and a standard deviation of 192. Find the probability that a randomly selected student who took the SAT-I has a greater score than 700. Round to 4 decimals
Answer #1

Solution :

Given that ,

mean = = 982

standard deviation = = 192

P(x > 700) = 1 - P(x < 700)

= 1 - P[(x - ) / < (700-982) /192 )

= 1 - P(z < -1.47)

= 1 - 0.0708 = 0.9292,Probability = 0.9292